# Lavida-O: Elastic Masked Diffusion Models for Unified Multimodal Understanding and Generation

### v1

@@ 3.1.2 Modality-aware Masking
3.2 Task-Specific Designs
Universal Text Conditioning. A common approach to improving the quality of text-to-image models is micro-conditioning (Bai et al., 2024; Podell et al., 2023), which conditions the image generation process on extra parameters such as original image resolution, crop coordinates, and image quality scores. This is typically achieved via specialized embeddings or modulation layers. However, since a unified model has strong language understanding and reasoning capabilities, we can simply append these conditions as plain text (e.g., “SCORE: 5.40”) to the end of user prompts. In addition to common conditions, we also incorporate image luminance and contrast as micro-conditions. This simple and effective design not only improves image quality by biasing generation toward high-scoring distributions, but also gives users more refined control over outputs. We provide additional details in Appendix A.4.
Stratified Random Sampling. Existing MDMs mostly adopt confidence-based sampling at inference, where tokens with high confidence scores are unmasked first. However, this design is suboptimal for image generation, where many tokens need to be decoded at each step. In particular, high-confidence tokens tend to cluster in small spatial regions adjacent to already unmasked tokens. This negatively affects image quality, since adjacent tokens are highly correlated, yet MDMs sample all unmasked tokens independently and assume their joint distribution perfectly factorizes. To mitigate this, we aim to minimize correlation between unmasked tokens at each step by maximizing their spatial distance. We achieve this with a stratified random sampling process, which ensures unmasked tokens are evenly distributed across the image. Specifically, we first divide the image into $2\times 2$ square regions and unmask one random token per region. We then subdivide each region into 4 subregions, giving 16 in total. Four already contain an unmasked token; we randomly unmask one token from each of the remaining 12. We repeat this process until all tokens are unmasked. This is illustrated in Figure 4(b). More details and analysis are provided in Appendix A.5 and B.3.
Planning and Reasoning. While existing unified MDMs integrate image understanding and generation tasks with a single diffusion objective, they do not incorporate mechanisms that use understanding to improve generation, except for the assumption that joint training benefits both tasks. To address this, we introduce two explicit mechanisms that leverage understanding to improve generation: planning and reflection. With planning, the model first generates a layout of the image represented by bounding boxes, then creates the actual image accordingly. For image editing tasks, it first identifies the desired edit region before generating the edited image. With reflection, the model evaluates its own generation using its understanding capability and determines whether it satisfies the user’s request. If misalignment is detected, the model generates a new image correcting the error. Examples are shown in Figure 1, with additional results in the appendix.
@@ 4.1 Setup
We start with LaViDa, an 8B understanding-only diffusion model, and extend it with a 2.4B image generation branch using the ElasticMoT architecture described in Section 3.1.1. The training consists of three stages:
  • Stage 1: We continue training the base model on object grounding and image-level understanding tasks using 12M samples. In this stage, the generation branch is not activated.
  • Stage 2: We pretrain the 2.4B image generation branch for text-to-image generation tasks using 200M image-text pairs. In this stage, we load 2.4B + 4B parameters using the ElasticMoT design described in Section 3.1.1. We freeze the understanding branch and train only the generation branch, thereby avoiding the need to include understanding data to prevent forgetting. To further improve training efficiency, we start at a resolution of 256 and progressively increase it to 512 and 1024 during training.
  • Stage 3: In the final stage, we jointly train the entire 2.4B + 8B model end-to-end on image-level understanding, text-to-image generation, image editing, and interleaved generation tasks such as planning and self-reflection. In addition to the data used in previous stages, we incorporate 1.5M image editing samples and 3M interleaved generation samples.
More details about the training data and process are provided in Apendix B.1.
4.2 Image Understanding
We report performance on image understanding tasks in Table 2. Lavida-O outperforms the previous state-of-the-art unified diffusion model, MMaDa, by a considerable margin on MMMU (Yue et al., 2024), MME (Fu et al., 2023), and MMB (Liu et al., 2024c). Compared with the base model LaViDa, Lavida-O achieves substantial improvements on most benchmarks such as ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021), ScienceQA (Lu et al., 2022), and MathVista (Lu et al., 2023), thanks to the scaling of training data.
Table 2: Quantitative results on image-level understanding tasks. *We evaluated results not reported in the original paper using official weights.
@@ 4.2 Image Understanding
4.3 Text-to-Image Generation
We report text-to-image generation results on the GenEval (Ghosh et al., 2023) and DPG (Hu et al., 2024) benchmarks, and FID scores on 30k prompts from MJHQ (Li et al., 2024b) dataset. We compare against text-to-image models including Flux-dev (Labs, 2024), SD3-Medium (Esser et al., 2024), Meissonic (Bai et al., 2024) and DALLE-3 (OpenAI, 2023), unified models including BAGEL (Deng et al., 2025), Show-o (Xie et al., 2024), MMaDa (Yang et al., 2025) and Muddit (Shi et al., 2025). Lavida-O significantly outperforms the state-of-the-art masked image generation model Meissonic, as well as unified models like MMaDa and Muddit. Planning and reflection further enhance prompt-following performance. We did not activate planning and reflection on MJHQ due to its large size and that FID does not reflect prompt-following capabilities.
Table 3: Quantative results on text-to-image generation tasks. *We evaluated results not reported in the original paper using official weights.
@@ 4.3 Text-to-Image Generation
4.4 Object Grounding
We evaluate the object grounding capabilities of Lavida-O on RefCOCO Referring Expression Comprehension (REC) tasks, reporting the Precision@0.5 metric. Lavida-O outperforms autoregressive vision-language models such as Qwen2.5-VL-7B (Bai et al., 2025) and InternVL3-8B (Zhu et al., 2025), as well as specialist models such as Grounding-DINO-L (Liu et al., 2024b) and SegLLM-7B (Wang et al., 2025a).
Table 4: Quantative results on RefCOCO, RefCOCO+, and RefCOCOg REC tasks. We report Precision@0.5 on object grounding tasks and compare against baseline models.
Model RefCOCO RefCOCO+ RefCOCOg
@@ 4.4 Object Grounding
LaViDa-O 92.3 94.8 89.0 88.7 92.5 83.3 90.0 90.6
4.5 Image Editing
We report image editing performance on the Image-Edit Bench in Table 5. Lavida-O outperforms state-of-the-art unified models such as BAGEL and specialized models like FluxKontext-dev. Most notably, Lavida-O even outperforms the state-of-the-art closed-sourced model GPT4-o(OpenAI, 2024) on replacing and removing objects, which requires localized understanding of the image. This underscores the effectiveness of Lavida-O’s design in integrating object grounding capabilities.
Table 5: Quantative results on Image-Edit benchmark. We report the per-category and overall score on Image-Edit Benchmark.
@@ 4.6 Training and Inference Speed
In Figure 5, we benchmark the inference efficiency of Lavida-O across three tasks: text-to-image generation, object grounding, and math reasoning. We measure end-to-end latency in seconds per image. Lavida-O is significantly faster than autoregressive models. Notably, we achieve a $6.8\times$ speedup on object grounding tasks compared to Qwen2.5-VL-7B (Bai et al., 2025). We also report the training efficiency measured by per-step latency and compare our Elastic-MoT design with BAGEL-style standard MoT design, Elastic-MoT improves the training speed by $3.17\times$. We provide additional analysis on the speed-quality tradeoff at inference time in Appendix B.6 and analysis on the training efficiency of Elastic-MoT design in B.2.
Figure 5: Training and Inference Speed of Lavida-O. We compare the end-to-end inference latency of Lavida-O on three distinct tasks: text-to-image generation, object grounding, and math reasoning. We also report the pretraining efficiency measured by per-step latency.
@@ 5 Conclusion
In summary, we proposed Lavida-O, the first multi-modal masked diffusion model that achieves state-of-the-art performance on text-to-image generation, image editing, and grounding tasks—competitive with the best specialist models and autoregressive unified models.
We also introduced a novel paradigm of interleaved generation, which explicitly leverages understanding capabilities to improve generation results in a unified multi-modal model through planning and self-reflection.
In developing Lavida-O, we proposed several efficient training and inference techniques, including the ElasticMoT architecture, universal text conditioning, and stratified random sampling, providing valuable insights for future work in masked diffusion models and unified multi-modal systems.

### v2

@@ 3.2 Task-Specific Designs
Stratified Random Sampling. Most MDMs use confidence-based sampling, unmasking high-confidence tokens first. In image generation, high-confidence tokens tend to cluster around already unmasked tokens. This negatively affecting image quality because adjacent tokens are highly correlated, which contradicts the independence assumption of MDMs. To mitigate this, we introudced a stratified sampling process. Starting with a $2\times 2$ grid, we unmask one token per region to ensure broad spatial coverage. Each region is then recursively subdivided into four smaller subregions, and we continue unmasking one token from each new region. This process repeats until all tokens are revealed, producing a balanced, evenly distributed unmasking pattern across the entire image. This is illustrated in Figure 4(b). More details and analysis are provided in Appendix A.5 and B.3.
Planning and Reasoning. While existing unified MDMs integrate image understanding and generation tasks with a single objective, they do not incorporate mechanisms that use understanding to improve generation, except for the assumption that joint training benefits both tasks. To address this, we introduce two explicit mechanisms that leverage understanding to improve generation: planning and reflection. With planning, the model first generates a layout of the image represented by bounding boxes, then creates the actual image accordingly. For image editing tasks, it first identifies the desired edit region before generating the edited image. With reflection, the model evaluates its own generation using its understanding capability and determines whether it satisfies the user’s request. If misalignment is detected, the model generates a new image correcting the error. Examples are shown in Figure 1, with additional technical results and analysis in Appendix A.7 and B.5.
@@ 4.1 Setup
We start with LaViDa (Li et al., 2025a) and extend it with a 2.4B image generation branch using the ElasticMoT architecture described in Section 3.1.1. The training consists of three stages: Stage 1: We continue training the base model on object grounding and image-level understanding tasks. Stage 2: We incroprate an 2.4B image generation and pretrain for text-to-image generation. We start with a resolution of 256 and progressively increase it to 512 and 1024 during training. Stage 3: In the final stage, we jointly train the entire 2.4B + 8B model end-to-end on image understanding, text-to-image generation, image editing, and interleaved generation tasks such as planning and self-reflection. More details on the training data and process are provided in Appendix B.1.
4.2 Main Results
Image Understanding. We report the performance of image understanding tasks in Table 2. Lavida-O outperforms the previous state-of-the-art unified diffusion model, MMaDa, by a considerable margin on MMMU (Yue et al., 2024), MME (Fu et al., 2023), and MMB (Liu et al., 2024c). Compared with the base model LaViDa, Lavida-O achieves substantial improvements on most benchmarks such as ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021), ScienceQA (Lu et al., 2022), and MathVista (Lu et al., 2023), due to the scaling of the training data.
Table 2: Quantitative results on image-level understanding tasks.*Evaluated by us.
@@ 4.2 Main Results
Text-to-Image Generation. We report text-to-image generation results on the GenEval (Ghosh et al., 2023) and DPG (Hu et al., 2024) benchmarks, and FID scores on 30k prompts from the MJHQ (Li et al., 2024b) dataset. We compare against text-to-image models including Flux-dev (Labs, 2024), SD3-Medium (Esser et al., 2024), Meissonic (Bai et al., 2024) and DALLE-3 (OpenAI, 2023), unified models such as BAGEL (Deng et al., 2025), MMaDa (Yang et al., 2025) and Muddit (Shi et al., 2025). Lavida-O significantly outperforms the state-of-the-art Meissonic masked image generation model, as well as unified models such as MMaDa and Muddit. Planning and reflection further enhance prompt-following performance. We did not activate planning and reflection on MJHQ due to its large size and that FID does not reflect prompt-following capabilities.
Table 3: Quantative results on text-to-image generation tasks. *Evaluated by us.
@@
Object Grounding. We evaluate the object grounding capabilities of Lavida-O on RefCOCO Referring Expression Comprehension (REC) tasks, reporting the Precision@0.5 metric. Lavida-O outperforms autoregressive vision-language models such as Qwen2.5-VL-7B (Bai et al., 2025) and InternVL3-8B (Zhu et al., 2025), as well as specialist models such as Grounding-DINO-L (Liu et al., 2024b) and SegLLM-7B (Wang et al., 2025a).
Table 4: Precision@0.5 on RefCOCO, RefCOCO+, and RefCOCOg REC tasks.
Model RefCOCO   $\uparrow$ RefCOCO+   $\uparrow$ RefCOCOg   $\uparrow$
@@
LaViDa-O (4-step) 92.3 94.8 89.0 88.7 92.5 83.3 90.0 90.6
LaViDa-O (1-step) 91.9 94.6 88.4 87.4 91.7 82.2 89.5 89.8
Image Editing. We report image editing results on Image-Edit Bench in Table 5. Lavida-O outperforms state-of-the-art unified models such as BAGEL and specialized models like FluxKontext-dev. Most notably, Lavida-O even outperforms the state-of-the-art closed-source model GPT4-o(OpenAI, 2024) on replacing and removing objects, which requires localized understanding. This underscores the effectiveness of Lavida-O’s design in integrating object-grounding capabilities.
Table 5: Per-Category and overall scores on Image-Edit benchmark.
@@ 4.3 Training and Inference Speed
In Figure 5, we benchmark the inference efficiency of Lavida-O across three tasks: text-to-image generation, object grounding, and math reasoning. We measure end-to-end latency in seconds per image. Lavida-O is significantly faster than autoregressive models. Notably, we achieve a $6.8\times$ speedup on object grounding tasks compared to Qwen2.5-VL-7B (Bai et al., 2025). We also report the training efficiency measured by per-step latency and compare our Elastic-MoT design with BAGEL-style standard MoT design, Elastic-MoT improves the training speed by $3.17\times$. Specifically, reducing the size of generation branch leads to a speedup of $2.23\times$, and decoupling the attention operation in the last 16 layers lead to an additional speedup of $1.44\times$, We provide additional analysis on the speed-quality tradeoff at inference time in Appendix B.6 and analysis on the training efficiency of Elastic-MoT design in B.2.
Figure 5: Training and Inference Speed of Lavida-O. We compare the end-to-end inference latency of Lavida-O on three tasks, as well as pretraining efficiency measured by per-step latency.
4.4 Additional Qualitative Results
@@ 5 Conclusion
In summary, we proposed Lavida-O, the first multi-modal masked diffusion model that achieves state-of-the-art performance on text-to-image generation, image editing, and grounding tasks—competitive with the best specialist models and autoregressive unified models. We also introduced a novel paradigm of interleaved generation, which explicitly leverages understanding capabilities to improve generation results in a unified multi-modal model through planning and self-reflection. In developing Lavida-O, we proposed several efficient training and inference techniques, including the ElasticMoT architecture, universal text conditioning, and stratified random sampling, providing valuable insights for future work in masked diffusion models and unified multi-modal systems.